OpenStudy (anonymous):

integrate log(sin^2x)

OpenStudy (xapproachesinfinity):

hmm any thoughts?

OpenStudy (xapproachesinfinity):

did you try anything

OpenStudy (xapproachesinfinity):

is that log base 10 or just ln?

OpenStudy (anonymous):

Its ln. My friend tried using integration by parts but it goes on infinitely.

OpenStudy (xapproachesinfinity):

hmm integration by part! what do you mean it goes to infinity do you there is no end the how many times you need to do integration by part

OpenStudy (xapproachesinfinity):

integration by part makes good start to me since u sub would not work

OpenStudy (anonymous):

true.

OpenStudy (xapproachesinfinity):

you might want to repeat that integration by part call your initial integral I=int (ln(sin^2x))dx

OpenStudy (anonymous):

ill try it out

OpenStudy (xapproachesinfinity):

yeah try and see if there is any cancellation otherwise the integration by part cannot solve this

OpenStudy (anonymous):

nope there isnt. it just gets more complicated.

OpenStudy (xapproachesinfinity):

hmm from wolfram seems there is something you can do http://www4b.wolframalpha.com/Calculate/MSP/MSP45861h30e7hbigg4ec5e00005972aaa8e0d8ac34?MSPStoreType=image/gif&s=56&w=521.&h=34.

OpenStudy (xapproachesinfinity):

that's a nasty solution by wolfram, but there should be a more interesting solution

OpenStudy (anonymous):

wow that was so complicated :/

OpenStudy (xapproachesinfinity):

hehe that's a crazy one! try you might find something interesting

OpenStudy (xapproachesinfinity):

i want to give it a try but i'm lazy hehehe

OpenStudy (anonymous):

or the questions prolly wrong :P

OpenStudy (xapproachesinfinity):

where did u get the question! i don't think it's wrong lol

OpenStudy (anonymous):

from my math book

OpenStudy (xapproachesinfinity):

what book?

OpenStudy (anonymous):

You prolly wont know the book

OpenStudy (xapproachesinfinity):

did you find: \[I=x\ln(\sin^2x)-2\int\limits x\cot x dx\]

OpenStudy (anonymous):

oh wait yeah. that was so stupid

OpenStudy (xapproachesinfinity):

just say the name of the book! lol is it stewart's book?

OpenStudy (xapproachesinfinity):

i think you can solve that second integral easily? why there is a problem

OpenStudy (anonymous):

i should have just expanded\[xsin2x/(1-\cos2x)\]

OpenStudy (anonymous):

lol i feel like an idiot.

OpenStudy (xapproachesinfinity):

expand what? why did you make your life hard

OpenStudy (xapproachesinfinity):

cotx is something that you can integrate

OpenStudy (anonymous):

yeah thanks

OpenStudy (xapproachesinfinity):

welcome

OpenStudy (xapproachesinfinity):

hmm i see you wrote 2sinxcosx=sin2x

OpenStudy (xapproachesinfinity):

darn the website is lagging again

OpenStudy (xapproachesinfinity):

a wise u v is u=x , dv=cotx

OpenStudy (xapproachesinfinity):

we know integral of cotx already

OpenStudy (anonymous):

yeah i did that and u end up with a bunch of stuff and integral log|sinx|

OpenStudy (xapproachesinfinity):

i found a nice solution

OpenStudy (xapproachesinfinity):

what's your result

OpenStudy (anonymous):

ugh i give up it keeps going on..... or im not in the mood to solve this. Show me the solution

OpenStudy (xapproachesinfinity):

hold on...

OpenStudy (xapproachesinfinity):

i made a mistake the result so far is \[\int\limits x\cot xdx=x\ln|\sin x|-\int\limits\ln(\sin x)dx\]

OpenStudy (anonymous):

yeah i got that

OpenStudy (xapproachesinfinity):

we need another integration by part this is a lot of fun carry the integration

OpenStudy (xapproachesinfinity):

i suppose that integral will be nice to deal with

OpenStudy (xapproachesinfinity):

you will probably go back to that cotx thingy

OpenStudy (anonymous):

i told u it goes on infinitely

OpenStudy (xapproachesinfinity):

no it won't believe me

OpenStudy (anonymous):

u ust said it urself

OpenStudy (anonymous):

it goes back to the cot thingy

OpenStudy (xapproachesinfinity):

you will have \[J=x\ln(\sin x)-\int\limits\ln(\sin x)dx=x\ln(\sin x)-x\ln(\sin x)+\int\limits x\cot xdx\]

OpenStudy (xapproachesinfinity):

where \[J=\int\limits x\cot xdx\]

OpenStudy (xapproachesinfinity):

nice cancellations in fact

OpenStudy (anonymous):

what wait we still have \[\int\limits(xcotx)dx\]

OpenStudy (xapproachesinfinity):

hmm there is a problem hehe i must of made some mistake somewhere

OpenStudy (xapproachesinfinity):

yes! i was hoping i would eliminate that j

OpenStudy (anonymous):

but that doesnt happen instead u have an INFINITE loop

OpenStudy (xapproachesinfinity):

that's not the question hehee

OpenStudy (anonymous):

bleh i dont think this problem is correct but if ur still determined that there is an answer then try solving it :P

OpenStudy (xapproachesinfinity):

yeah, I'm still sure that it ends somewhere buy you still don't want to tell me what's name of the book lol

OpenStudy (anonymous):

its from a sample paper for some competetive exam

OpenStudy (xapproachesinfinity):

well name them lol! you don't have to anyway, hmm... i see that's why it a little of nasty but it is not all complicated!

OpenStudy (anonymous):

haha solve it and show me the final answer then

OpenStudy (xapproachesinfinity):

i still want to try one more trick hehe but time is running out lol thinking about substitution for \[\int\limits \ln(\sin x)dx\]

OpenStudy (xapproachesinfinity):

if i didn't solve it, i will come back with the solution later once i have time don't close the question

OpenStudy (anonymous):

k sure.

OpenStudy (xapproachesinfinity):

ideas are floating hehe damned looping integral, i will solve it no matter what

OpenStudy (xapproachesinfinity):

seems my way was wrong hehe

OpenStudy (xapproachesinfinity):

@ganeshie8

OpenStudy (anonymous):

hey forget it. My friend was being stupid and forgot to tell me that it was a definite integrals problem That makes the whole thing much easier

OpenStudy (xapproachesinfinity):

yeah what are the limits

OpenStudy (xapproachesinfinity):

it should have some solution even if it's indefinite

OpenStudy (xapproachesinfinity):

i got stuck with \[\int\limits \ln(\sin x)dx\] integration by part doesn't work

OpenStudy (xapproachesinfinity):

i had enough of this one lol, time to sleep hehe